OpenStudy (anonymous):

The length of a rectangle is 2 meters more than its width. If the area is 35 square meters, what are the length and width? A 20-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 4 ft longer than the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace. Please help I'm absolutely lost on these two. Thank You.

OpenStudy (anonymous):

x=length y=width xy=35 x=y+2

OpenStudy (anonymous):

(y+2)(y)=35 y^2+2y=35 y^2+2y-35=0 (y-5)(y+7) y = 5 or -7 however y cannot be - so y = 5 x=5+2 = 7 y=5 x=7

OpenStudy (anonymous):

do you have a picture of the second question

OpenStudy (anonymous):

yes i believe a picture would be nice

OpenStudy (anonymous):

they don't give me a picture for the second one

OpenStudy (anonymous):

just the question to answer that's why its kind of confusing for me

OpenStudy (anonymous):